OpenStudy (anonymous):

3y=x+6. What is the slope?

OpenStudy (anonymous):

You want to get your equation into slope intercept form y=mx+b Once you do that your slope will be m

OpenStudy (anonymous):

It looks like some of your post is cut off. Can you please re-post the ending starting at slope will be? thanks

OpenStudy (anonymous):

sorry comp. kicked me off

OpenStudy (anonymous):

Ok so you want to get y by itself so that your equation is in slope incercept form: y=m(x)+b Your slope is m

OpenStudy (anonymous):

So for this specific equation you want to divide each side by three to get y alone y=1/3x+1/3b your slope should be 1/3 Make sense?

OpenStudy (anonymous):

ok so I get .33333333 right?

OpenStudy (anonymous):

Yes 1/3 in decimal form

OpenStudy (anonymous):

Is there another place to get help. This site seems to lock up a lot.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Not that I have found.... it seems like it has been really slow today and no problem!

OpenStudy (anonymous):

it has locked up on me several times and i had to refresh each time.

OpenStudy (anonymous):

Ya Idk what the deal is...